Lincoln High School's basketball team won the regional playoffs scoring a total of 60 points not including free throws. The team made a total of 26 basket; some were 2-points shots and the rest were 3-points shots. How many 2-points shots did the team make?
Answer by rothauserc(4718)

let x be the number of 2 point baskets made and y be the number of 3 point baskets made
:
1) x + y = 26
:
2) 2x + 3y = 60
:
solve equation 1 for x and substitute in equation 2
:
x = 26 - y
:
2(26 - y) +3y = 60
:
52 -2y +3y = 60
:
y = 8
:
x = 26 - 8 = 18
:
******************************
The team made 18 2-point shots
******************************
: